OpenStudy (anonymous):

A university surveys 36 randomly selected students and asks the number of dollars spent on textbooks/supplies during the academic year. The results in dollars are shown below: 425 78 53 699 663 282 845 373 161 22 853 653 366 806 920 562 828 931 641 134 916 38 379 90 388 894 654 639 916 284 700 684 600 425 717 822 With a 90% confidence interval. would a given mean value of $600 be likely given the calculated confidence interval? 90% confidence interval (_______________=/-_____________)

OpenStudy (anonymous):

After computing the mean value of the dataset , you must evaluate the allowed boundaries of the 90% confidence interval (wich is in fact that the value must be 100-90=10% up or down). If the value fall inside the 10%less-10%more interval, it is in the specified confidence interval.

OpenStudy (anonymous):

Need more details please! I got 459.15 and 620.86

OpenStudy (anonymous):

It seems to me that you computed the 90% interval out of the given 600. However, you actually have to find the real mean of your dataset (about 540), and then do 540 - 10% = 486, and 540+10% = 594. Is 600 between 486 and 594 ? No, it isn't : therefore, the given mean value is incorrect in regard of the given confidence interval 486/594.